# NovelQA: Benchmarking Question Answering on Documents Exceeding 200K Tokens

### v2

@@ E.1 Implementations
We also tested Gemini-1.5555https://deepmind.google/technologies/gemini/flash/, but generation for 816 questions from 29 novels was blocked for unknown reasons (with only 14 of these novels under copyright protection). Consequently, we have decided not to present Gemini’s result.
E.2 Multi-choice Performance on Each Type of question
@@
Besides the evaluation in generative settings, we also prompted the models to collect their responses for the multi-choice versioned questions. Table 11 presents the accuracies of four models in multi-choice settings in each question type.
E.3 Relation between Multi-choice Accuracy and Evidence Positions
@@
Figure 6: Analysis of Accuracy in multi-choice Setting by Absolute and Relative Token Positions: This figure illustrates the accuracy in the multi-choice setting of NovelQA, plotted against the absolute position (left) and percentage position (right) of each question’s relevant evidence within the novel. Each subplot represents a different model. The x-axis of the absolute position figure (left), reflecting token indexes, is folded on the right due to the long-tail distribution in the lengths of the selected novels.
@@
Figure 6 which presents the relationships between the accuracy and the absolute or relative positions accordingly shows similar trends to which are observed in the generative setting. To be specific, the accuracy by absolute token position remains high when related evidences are before 100K’s text length, while drops after 100K. Meanwhile, the accuracy by relative position remains relatively even. A comparison is made between the accuracies within two ranges, 65K(the lowest token count) to 100K (namely pre-100K) and 100K to the end (namely post-100K). Figure 7 and Table 10 present a clearer contrast between these two ranges, where the accuracies drop sharply after the 100K token.
@@
Figure 7: Performance Analysis by Relative Positions within Different Token Ranges: This figure illustrates the accuracy of various long-context LLMs in the multi-choice setting, segmented by novels’ length categories: 65K(the lowest token count) to 100K tokens and over 100K tokens. It highlights the models’ performance trends relative to the evidence position within these ranges, showcasing the ‘lost-in-middle’ phenomenon in shorter texts and a performance decline towards the end in longer texts, thus revealing distinct comprehension challenges faced by LLMs when processing texts of varying lengths.
E.4 Representative Errors
@@
Through reviewing the generation above, we concluded three common error types: hallucination, miscounting and overlooking. Examples for each type are presented in Table 12.
@@
Overlooking refers to the model’s neglect of details. As mentioned above, the questions in detailed category involve minor characters, plots, or settings. Diving further, the reasons why these details are difficult to be recalled lie in two aspects: (1) They do not contribute to the character development, other plots, or the main themes, and thus reading the rest of the novel does not help to remind this detail; (2) Since most novels have derivative works (e.g., films, fan works, and book reviews), where the detailed information is eliminated to form a condensed narration. As the derivative works spread further and appear more frequently in the model’s training data, they have a higher probability of becoming the models’ inner knowledge, which is similar to [Chang et al., 2023]’s observation, and vice versa for those omitted details. These two factors contribute to the difficulty in the model’s recalling details and thus result in overlooking errors.
Miscounting Researches [Li et al., 2023a, Feng et al., 2023] has revealed shortcomings in the counting ability of LLMs, especially autoregressive-decoder-based models, and methods unfolding the outputs such as chain-of-thought prompting can enhance their counting ability. Our test does show that models make mistakes with numbers. Though the errors in the case of generative responses may be due to not following instructions and simply outputting ‘multiple times’ instead of the desired specific times, the accuracy in the multi-choice setting has still only reached 38.53% to 49.56% for the chosen four models, as shown in Table . Even in the simplest question which asks for the appearing frequency of certain phrases, the models still make mistakes.
Close-book QA
Multichoice Generative
GPT-4 60.94 (-9.06) 34.30 (-13.58)
Claude 2.1 51.77 (-15.01) 22.36 (-20.68)
InternLM2-Chat-7b 33.58 (-10.71) 14.12 (-16.81)
InternLM2-Chat-20b 33.05 (-16.13) 15.51 (-16.86)
E.5 Close-book Question Answering
We employ a Close-book QA scenario to assess the extent to which models rely on the content of novels versus using their internal knowledge to answer questions. In this approach, the models are not given access to the text of the novels and must rely solely on their pre-existing knowledge to provide answers. Given that our selected novels are well-known and representative of their genres, it’s inevitable that LLMs have encountered their texts during training and retained some of their content. The results of this evaluation are presented in Table 13.
Models like GPT-4 and Claude 2.1 achieve notable scores in the Close-book setting (60.94% and 51.77% in multi-choice, 34.30% and 22.36% in generative, respectively) indicating that they have internalized significant portions of the novels’ content during training. This internal knowledge allows them to perform reasonably well even without direct access to the text. The difference in performance between the Close-book and standard settings underscores the challenges in long-context understanding. While models can retain and recall information from well-known texts, their ability to comprehend and use such information to answer questions accurately diminishes in the absence of the text. This suggests that long-context understanding, as measured by the main results, might still be more challenging than it appears, as models benefit from having the text directly available.

### v3

@@ C.4.1 Implementations
We also tried to test Gemini-1.5444https://deepmind.google/technologies/gemini/flash/, but the generation of 816 questions from 29 novels was blocked for unknown reasons (with only 14 of these novels under copyright protection). Consequently, we have decided not to present Gemini’s result.
Experiment Configs Given the cost of running long-context APIs, we request that each model respond to all questions for a given book in a single session. To ensure fair comparisons, local-deployed LLMs also answer all questions for a book at once. We set ‘temperature = 0’ to eliminate randomness and keep other hyper-parameters default.
C.4.2 Evaluator Model
Table 15: Two evaluator models’ (GPT-4o-mini and Claude-3.5-sonnet) evaluations on the answers of these two models. The results show minimal variance between evaluators, with Claude-3.5 being slightly stricter but showing no significant model preference.
Evaluated By GPT-4o-mini Evaluated By Claude-3.5-sonnet
Claude-3.5-sonnet 62.30 61.39
GPT-4o-mini 52.32 51.56
To prove that the model bias has little or no effect on our final results, we conducted a thorough analysis comparing different evaluator models to check their potential, demonstrated in Table 15. The results show minimal variance between evaluators, with Claude-3.5 being slightly stricter but showing no significant model preference. This consistency is largely due to NovelQA’s objective nature, with well-defined questions and answers that leave little room for evaluator bias.
C.4.3 Multichoice Performance Related to Question Type
@@
Besides the evaluation in generative settings, we also prompted the models to collect their responses for the multichoice versioned questions. Table 14 presents the accuracies of four models in multichoice settings in each question type.
C.4.4 Multichoice Performance Related to Evidence Position
@@
Figure 6: Analysis of Accuracy in multichoice Setting by Absolute and Relative Token Positions: This figure illustrates the accuracy in the multichoice setting of NovelQA, plotted against the absolute position (left) and percentage position (right) of each question’s relevant evidence within the novel. Each subplot represents a different model. The x-axis of the absolute position figure (left), reflecting token indexes, is folded on the right due to the long-tail distribution in the lengths of the selected novels.
@@
Figure 6 which presents the relationships between the accuracy and the absolute or relative positions accordingly shows similar trends to those observed in the generative setting. To be specific, the accuracy by absolute token position remains high when related evidences are before 100K’s text length, while it drops after 100K. Meanwhile, the accuracy by relative position remains relatively even. A comparison is made between the accuracies within two ranges, 65K(the lowest token count) to 100K (namely pre-100K) and 100K to the end (namely post-100K). Figure 7 and Table 4 present a clearer contrast between these two ranges, where the precision drops dramatically after the 100K token.
@@
Figure 7: Performance Analysis by Relative Positions within Different Token Ranges: This figure illustrates the accuracy of various long-context LLMs in the multichoice setting, segmented by novels’ length categories: 65K(the lowest token count) to 100K tokens and over 100K tokens. It highlights the models’ performance trends relative to the evidence position within these ranges, showcasing the ‘lost-in-middle’ phenomenon in shorter texts and a performance decline towards the end in longer texts, thus revealing distinct comprehension challenges faced by LLMs when processing texts of varying lengths.
C.4.5 Multi-hop Accuracy Related to Evidence Distance
Table 16: Correlation between the distance among multiple evidences in multi-hop questions and the accuracy in the generative scenario. The max distance among all evidence distances are considered as the evidence distance of the question. The overall negative correlations show that the evidence distances are negatively correlated with the accuracies. The interpretation of matrices can be found in Appendix C.4.5.
Model Pearson Spearson Kendall
GPT-4 -0.0734 (0.0879) -0.2511 (3.0696×$\times e^{-09}$) -0.2031 (4.6129×$\times e^{-09}$)
Claude-3 -0.0120 (0.7899) -0.1826 (4.6228) -0.1385 (9.0407×$\times e^{-05}$)
InternLM-7b -0.1140 (0.0083) -0.2287 (8.8721×$\times e^{-08}$) -0.1841 (1.2663×$\times e^{-07}$)
InternLM-20b -0.0634 (0.1407) -0.2017 (2.1954×$\times e^{-06}$) -0.1626 (2.722×$\times e^{-06}$)
We also measured the relationship between the evidence distance within each multi-hop question and the accuracy under the generative task. For each multi-hop question, we obtain the distances among all evidences, and consider the max distance among them as the evidence distance. This can be interpreted as the model must memorize at least one of its evidences for the max distance to meet the final evidence in order to obtain the answer. The correlation between the evidence distance and the accuracy for multi-hop questions is demonstrated in Table 16. Among the indices, Pearson correlation assumes that the two input distributions have linear correlation. Spearman correlation assumes that the two input distributions have monotonic correlation. Kendall correlation assumes that the two input distributions have ordinal correlation (in ranks). The results suggest that the max distance among evidence has a negative correlation between the accuracy and the max distance among evidences, which can be interpreted as a lower accuracy is expected on the questions with more distantly distributed evidences.
C.4.6 Representative Errors
@@
Through reviewing the generation above, we concluded three common error types: hallucination, miscounting, and overlooking. Examples for each type are presented in Table 17.
@@
Overlooking refers to the model’s neglect of details. As mentioned above, the questions in detailed category involve minor characters, plots, or settings. Diving further, the reasons why these details are difficult to be recalled lie in two aspects: (1) They do not contribute to the character development, other plots, or the main themes, and thus reading the rest of the novel does not help to remind this detail; (2) Since most novels have derivative works (e.g., films, fan works, and book reviews), where the detailed information is eliminated to form a condensed narration. As the derivative works spread further and appear more frequently in the model’s training data, they have a higher probability of becoming the models’ inner knowledge, which is similar to (Chang et al., 2023)’s observation, and vice versa for those omitted details. These two factors contribute to the difficulty in the model’s recalling details and thus result in overlooking errors.
Miscounting Researchers (Li et al., 2023a; Feng et al., 2023) have revealed shortcomings in the counting ability of LLMs, especially autoregressive-decoder-based models, and methods unfolding the outputs such as chain-of-thought prompting can enhance their counting ability. Our test does show that models make mistakes with numbers. Though the errors in the case of generative responses may be due to not following instructions and simply outputting ‘multiple times’ instead of the desired specific times, the accuracy in the multichoice setting has still only reached 38.53% to 49.56% for the chosen four models, as shown in Table 17. Even in the simplest question, which asks for the appearing frequency of certain phrases, the models still make mistakes.